Alfalfa meal contains 2% nitrogen by weight. Bat guano contains 10% nitrogen by weight. How many pounds of each should be mixed together to make 7 pounds of fertilizer that is 8% nitrogen?
Answer by stanbon(75887)

Alfalfa meal contains 2% nitrogen by weight. Bat guano contains 10% nitrogen by weight. How many pounds of each should be mixed together to make 7 pounds of fertilizer that is 8% nitrogen?
--------------------------------------
Let "A" be the weight of Alfalfa meal.
Let "B" be the weight of Bat guano.
--------------------------------------------
Equation:
Weight Equation: A + B = 7
nitrogen + nitrogen = nitrogen
0.02A + 0.10(7-A) = 0.08(7)
---
Multiply thru by 100 and solve for "A":
2A + 10*7 - 10A = 8*7
-8A = -3*7
A = 3/8 lbs (amt of Alfalfa meal needed in the mixture)
----
B = 7-(3/8) = (56/8)-(3/8) = 53/8 lbs (amt of Bat guano needed in the mix)
=============================================
Cheers,
Stan H.